### v2

# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers. Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer. The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities. However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs). KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs. In this work, we compare the performance of KANs against that of MLPs in graph learning tasks. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks. Code is available at https://github.com/RomanBresson/KAGNN.
1 Introduction
Graphs are structural representations of information which are useful for modeling many types of data. They arise naturally in a wide range of application domains, and their abstract nature offers increased flexibility. Typically, the nodes of a graph represent entities, while the edges capture the interactions between them. For instance, in social networks, nodes represent individuals, and edges represent their social interactions. In chemo-informatics, molecules are commonly modeled as graphs, with nodes corresponding to atoms and edges to chemical bonds. In other settings, molecules can also be nodes, with edges capturing their ability to bond with one another.
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [12]. Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [36]. Most GNNs consist of a series of message passing layers. Within a message passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
A lot of recent work has focused on investigating the expressive power of GNNs [22]. There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [27]. Furthermore, an equivalence has also been established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [3]. This line of work gave insights into the limitations of different models [38, 29], but also led to the development of more powerful architectures [27, 24, 26].
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [5, 15]. The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Having said that, in practice the models suffer from several limitations due to non-convex loss functions, algorithms without convergence guarantees and a notorious lack of interpretability that hinders their applicability in several domains. Recently, Kolmogorov-Arnold Networks (KANs) [23] have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem [21] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs, with learnable activations based on B-splines and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions, while simultaneously being more interpretable.
In this paper, we present a thorough empirical comparison between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our work is orthogonal to prior work that studies the expressive power of GNNs. Here, we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the model’s performance. We evaluate the different GNN models on several standard node classification, graph classification and graph regression datasets.
The rest of this paper is organized as follows. Section 2, provides an overview of the tasks we address in this paper, as well as a description of message passing GNNs and Kolmogorov-Arnold networks. In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In Section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks. Finally, Section 5 concludes the paper.
2 Background
2.1 Considered Graph Learning Tasks
Before presenting the tasks on which we focus in this study, we start by introducing some key notation for graphs. Let $\mathbb{N}$ denote the set of natural numbers, i. e., $\{1,2,\ldots\}$. Then, $[n]=\{1,\ldots,n\}\subset\mathbb{N}$ for $n\geq 1$. Let $G=(V,E)$ be an undirected graph, where $V$ is the vertex set and $E$ is the edge set. We denote by $n$ the number of vertices and by $m$ the number of edges, i. e., $n=|V|$ and $m=|E|$. Let $g\colon V\rightarrow[n]$ denote a bijective mapping from the space of nodes to set $[n]$. Let $\mathcal{N}(v)$ denote the the neighbourhood of vertex $v$, i. e., the set $\{u\mid\{v,u\}\in E\}$. The degree of a vertex $v$ is $\deg(v)=|\mathcal{N}(v)|$. Each node $v\in V$ is associated with a $d$-dimensional feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$. Thus, $\mathbf{x}_{v}$ is equal to the $g(v)$-th row of $\mathbf{X}$.
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn a function that maps nodes to their class labels, i. e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., to learn a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=\texttt{EMBEDDING}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=\texttt{READOUT}(\mathbf{H})$ for the entire graph.
One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes . Indeed, these orderings do not hold any semantic meaning and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max or mean operators.
2.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message passing framework [12] which consists of a sequence of layers and whithin each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
$\mathbf{h}_{v}^{(\ell)}=\phi^{(l)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
where $\bigoplus$ is a permutation-invariant aggregation function (e. g., mean, sum), and $\phi^{(\ell)}$ is a differentiable function (e. g., linear transformation, MLP) that combines and transforms the node’s previous embedding with the aggregated vector of its neighbors.
As discussed above, in this paper, we focus on the functions that different GNN models employ to update node representations. Many existing GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [7, 20, 40]. For instance, each layer of the the Graph Convolutional Network (GCN) [20] is defined as follows:
$\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (2)
where $\sigma$ is a non-linear activation and $\mathbf{W}^{(\ell)}$ is a trainable weight matrix.
However, the 1-layer perceptron is not a universal approximator of multiset functions [38]. Thus, the emerging GNN might not be expressive enough for some tasks. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [2, 6, 28, 29]. It is well-known that standard message passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [38]. While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) [38], which is defined as follows:
$\mathbf{h}^{(\ell)}_{v}=\texttt{MLP}^{(\ell)}\left((1+\epsilon^{(\ell)})\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (3)
where $\epsilon^{(\ell)}$ denotes a trainable parameter, and $\texttt{MLP}^{(\ell)}$ a trainable MLP.
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned. However, in practice, GIN might fail to learn those weights due to limited training data and due to limitations of the employed training algorithm (e. g., stochastic gradient descent). This has motivated a series of works which focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [13]. Other works have studied how to initialize the weights of the MLPs of the message passing layers of GNNs. It was shown that by adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [14]. On the other hand, there exist settings where there is no need of complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [35] or by feeding the node features into a neural network which generates predictions and then propagate those predictions via a personalized PageRank scheme [9].
2.3 Kolmogorov-Arnold Networks
Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community [23]. As mentioned above, this model relies on the Kolmogorov-Arnold representation theorem, which states that any multivariate function $f:\left[0,1\right]^{d}\rightarrow\mathbb{R}$ can be written as:
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\Phi_{i}\left(\sum_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})\right)$ (4)
where all $\Phi_{\square}$ and $\phi_{\square}$ functions are univariate, and the sum is the only multivariate operator.
Equation (4) can be seen as a two-step process. First, a different set of univariate non-linear activation functions is applied to each dimension of the input, and then the output of those functions are summed up. The authors rely on this interpretation to define a Kolmogorov-Arnold Network (KAN) layer, which is a mapping between a space $A\subseteq\mathbb{R}^{d}$ and a different space $B\subseteq\mathbb{R}^{d^{\prime}}$, (identical in use to an MLP layer). Such a layer consists of $d\times d^{\prime}$ trainable functions $\{\phi_{ij},~{}1\leq i\leq d^{\prime},~{}1\leq j\leq d\}$. Then, for $\mathbf{x}\in A$, we compute its image $\mathbf{x}^{\prime}$ as:
$\mathbf{x}_{i}^{\prime}=\sum\limits_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})$ (5)
Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation (4), and the derived model is a universal function approximator.
This seemingly offers a complexity advantage compared to MLPs, since the number of univariate functions required to represent any multivariate function from $\left[0,1\right]^{d}$ to $\mathbb{R}^{d^{\prime}}$ is at most $(2d^{2}+d)\times d^{\prime}$, whereas the universal approximation theorem for the MLP requires a possibly infinite number of neurons. However, as stated by the original paper, the behavior of such univariate functions might be arbitrarily complex (e. g., fractal, non-smooth), thus leading to them being non-representable, and non-learnable.
MLPs relax the arbitrary-width constraint by stacking finite-width layers. Likewise, KANs relax the arbitrary-complexity constraints on the non-linearities by stacking KAN layers. Thus, the output of a function is given by:
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})$ (6)
where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.
The original paper uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture.
Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [4] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [31]. So far, most efforts have focused on time series data [39]. For instance, KANs have been evaluated in the satellite traffic forecasting task [34]. Furthermore, they have been combined with architectures that are traditionally leveraged in time series forecasting tasks such as the Long Short-Term Memory Network [11] and the Transformer [10]. The work closest to ours is the one reported in [37], where the authors propose FourierKAN-GCF. This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.
3 KAN-based GNN Layers
We next derive variants of the GIN and GCN models which use KANs to transform the node features instead of fully-connected layers or MLPs.
3.1 The KAGIN Layer
To achieve its maximal expressivity, the GIN model relies on the MLP architecture and its universal approximator property. Since KAN is also a universal function approximator, we could achieve the same expressive power using KANs in lieu of MLPs. We thus propose the KAGIN model which is defined as follows:
$\mathbf{h}^{(\ell)}_{v}=\texttt{KAN}^{(l)}\left((1+\epsilon)\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (7)
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with arbitrary layer width. While this is not guaranteed with the spline-based implementation with limited grid size, the empirical results in the original paper demonstrate the great expressive power of KANs [23], especially of small models and setting where regularity is expected.
3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
$\mathbf{h}^{(\ell)}_{v}=\Phi^{(\ell)}\left(\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (8)
where $\Phi^{(\ell)}$ denotes a single KAN layer. In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as:
$\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ (9)
where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph.
4 Empirical Evaluation
In this section, we compare the KAGIN and KAGCN models with the GIN and GCN models in the following tasks: node classification, graph classification and graph regression. The code for reproducing the results is available at https://github.com/RomanBresson/KAGNN. All models are implemented with PyTorch [30]. For KAN layers, we rely on a publicly available implementation [^1] https://github.com/Blealtan/efficient-kan.
4.1 Node classification
Datasets.
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [20] and Ogbn-arxiv [16]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks. The homophilic networks are already split into training, validation and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross validation indices.
Experimental setup.
For every dataset and model, we tune the values of the hyperparameters. Specifically, we choose the values that perform best on the validation set (i. e., lowest validation error). To find these values, we use the Optuna package [1]. We set the number of iterations of Optuna equal to $100$ trials.
For all models, the learning rate is chosen from $[10^{-3},10^{-2}]$, the number of message passing layers from $\{1,2,3,4\}$, the hidden dimension size from $\{8,9,\ldots,128\}$ and the weight decay from $\{0,0.0005\}$. For KAGIN and KAGCN, we also choose the grid size from $\{3,4,5\}$ and the spline order from $\{1,2,3,4\}$. Once the best hyperparameter values are found, we evaluate the models on the test set. For the homophilic networks, we initialize and train $10$ different models (we use $10$ different random seeds). We then evaluate the $10$ models on the test set and report the average accuracy. For the heterophilic datasets, we tune each model’s hyperparameters within each fold and we report the average accuracy across the $10$ folds.
Table 1: Average classification accuracy ($\pm$ standard deviation) of the KAGIN, GIN, KAGCN, and GCN models on the $7$ node classification datasets.
Dataset KAGIN GIN KAGCN GCN
Cora 0.7620 $\pm$ 0.0077 0.5528 $\pm$ 0.1112
0.7826
$\pm$ 0.0177
0.7566 $\pm$ 0.0718
Citeseer 0.6837 $\pm$ 0.0117 0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185
0.6891
$\pm$ 0.0074
Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2110 $\pm$ 0.0628
0.2997
$\pm$ 0.0274
0.2777$\pm$ 0.0699
Cornell
0.5243
$\pm$ 0.0813
$0$$\pm$ 0.1414 0.4892 $\pm$ 0.0547 0.4162 $\pm$ 0.0737
Texas 0.5973 $\pm$ 0.0505
0.6054
$\pm$ 0.0516
0.5811 $\pm$ 0.0440 0.5892 $\pm$ 0.0415
Wisconsin
0.5510
$\pm$ 0.0720
0.4961 $\pm$ 0.1532 0.5196 $\pm$ 0.1020 0.5216 $\pm$ 0.0729
Actor
0.2890
$\pm$ 0.0139
0.2817 $\pm$ 0.0110 0.2755 $\pm$ 0.0134 0.2852 $\pm$ 0.0129
Results.
The results are given in Table 1. We can see that KAGIN outperforms GIN on all but one datasets. On some datasets, the difference in performance between the two models is significant. For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $20.9\%$ in accuracy over GIN. On the other hand, KAGCN outperforms GCN on $3/7$ considered datasets, albeit being the best in $2/3$ homophilic networks and exhibiting very similar accuracy in $3/4$ heterophilic ones. The KAGIN and GIN models exhibit the best performance in heterophilic networks while the KAGCN and GCN models achieve the highest accuracy on the homophilic. The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals, an overlap that is more prevalent with GCN-based models in $6/7$. Overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture. Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.g. GCN accuracy with heterophilic networks [41], however, it improves the models in the majority of the results i.e. $5/7$.
With regards to the optimal hyperparameters, we found that KAGIN required a larger grid size on average compared to KAGCN. In general, a grid size of $4$ was chosen in $8/14$ experiments and the predominant spline order was $1$, appearing in $8/14$ experiments while $3$ appears $2$ times and $4$ only once. The number of hidden layers and their sizes varied significantly through datasets but overall KAGIN had a substantially larger average hidden layer size i. e., $75.4$ compared to $48$ for KAGIN, as $128$ was chosen for $3$ datasets compared to only $1$ for KAGCN. This is in contrast to previous findings contending that GIN-based models requires more complex learning procedures, a pattern that seems to withstand with KANs.
Training times.
We present in Table 2 the training time per epoch for different configurations of the KAGIN and GIN models. We observe that for a given number of message passing layers and hidden dimension size, the KAGIN model is computationally more expensive than GIN. If the grid size and spline order hyperparameters of KAGIN are set to $1$, the difference in running time between the two models is very small (less than $0.02$ seconds per epoch for all configurations). However, the complexity of KAGIN increases are the grid size and spline order increase. Overall, our results suggest that the running time of KAGIN is slightly greater than that of GIN, and by no means prohibitive.
Table 2: Training time per epoch on the Ogbn-arxiv dataset.
Architecture
Message Passing
Layers
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 2 16 NA NA 5,336 0.035
GIN 2 64 NA NA 27,368 0.041
GIN 4 16 NA NA 6,936 0.037
GIN 4 64 NA NA 52,200 0.056
KAGIN 2 16 1 1 31,232 0.046
KAGIN 2 64 1 1 63,488 0.048
KAGIN 2 16 4 1 54,656 0.059
KAGIN 2 16 4 3 70,272 0.146
KAGIN 4 16 1 1 38,400 0.056
KAGIN 4 64 1 1 116,736 0.072
KAGIN 4 64 4 1 204,288 0.109
KAGIN 4 64 4 3 262,656 0.298
4.2 Graph Classification
Datasets.
In this set of experiments, we compare the KAGIN model against GIN on standard graph classification benchmark datasets [25]. We experiment with the 7 following datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M. The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
Experimental setup. We follow the experimental protocol proposed in [8]. Thus, we perform $10$-fold cross-validation to obtain an estimate of the generalization performance of each method, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the splits provided in [8]. We use the Optuna package to select the model that achieves the lowest validation error. We set the number of iterations of Optuna equal to $100$.
For a fair comparison, we set the number of message passing layers of both models to a fixed value for each dataset. Based on preliminary experiments, on MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. To produce graph representations, we use the sum operator. The produced graph representations are fed to an MLP (for GIN) or a KAN (for KAGIN) layer which computes the output. We train each model for $1,000$ epochs by minimizing the cross entropy loss. We use the Adam optimizer for model training [19]. We apply batch normalization [17] and dropout [33] to the output of each message passing layer. We also use early stopping with a patience of $20$ epochs. For both models, we choose the number of hidden layers from $\{2,3,4\}$ and the dropout rate from $[0.0,0.5]$. For GIN, we choose the hidden dimension size from $\{8,9,\ldots,256\}$ and the learning rate from $[10^{-5},10^{-2}]$. For KAGIN, we choose the hidden dimension size from $\{2,3,\ldots,128\}$, the grid size from $\{1,2,\ldots,16\}$, the spline order from $\{1,2,\ldots,8\}$ and the learning rate from $[10^{-4},10^{-2}]$. For the IMDB-B and IMDB-M datasets, where there are no node features, we annotate nodes with one-hot encodings of their degrees, up to $35$ (all degrees above $35$ are set equal to $35$). Once the best hyperparemeters are found for a given split, we train $3$ different models on the training set of the split (in order to limit the impact of random initialization) and evaluate them on the test set of the split. This yields $3$ test accuracies, and we compute their average.
Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $6$ graph classification datasets.
MUTAG DD NCI1 PROTEINS
GIN 85.09 $\pm$ 5.97 73.68 $\pm$ 3.89 79.57 $\pm$ 1.27 70.26 $\pm$ 3.62
KAGIN
85.45
$\pm$ 7.38
75.46
$\pm$ 4.16
79.72
$\pm$ 1.79
71.78
$\pm$ 3.12
ENZYMES IMDB-B IMDB-M
GIN
53.72
$\pm$ 9.23
73.03 $\pm$ 2.86
49.78
$\pm$ 3.38
KAGIN 42.94 $\pm$ 5.69
73.83
$\pm$ 4.35
49.78
$\pm$ 4.69
Results.
Table 3 illustrates the average classification accuracies and the corresponding standard deviations of the two models on the different datasets. We observe that the two models achieve similar levels of performance. KAGIN outperforms GIN on $5$ out of the $7$ datasets. However, the difference in performance between the two models is very small. This suggests that the two models are similar in terms of expressive power. On ENZYMES, however, KAGIN was found to perform much worse than GIN. Note that ENZYMES consists of more classes ($6$ classes in total) than the rest of the datasets, while ENZYMES and PROTEINS are the only datasets where the nodes of the graphs are annotated with continuous features (on the rest of the datasets, they are annotated with one-hot encodings). We hypothesize that the difference in performance is due to the inability of KANs to handle those continuous features. We thus normalized the node features within each fold by removing the mean of each feature (computed from the training samples) and then dividing by the corresponding standard deviation. We re-conducted the experiment and the average accuracy increased by approximately $6\%$ ($48.77\%$ instead of $42.94\%$). Therefore, it turns out that in some settings it might be harder for KAN layers to handle continuous features than for MLPs.
Training times.
We give in Table 4 an overview of the training times for different GIN and KAGIN architectures. We provide the number of parameters of each architecture. We notice that, for the same number of parameters, KAN is slower than its MLP counterpart. This is particularly sensitive to grid size and spline order. This makes sense since, using splines, each parameter involves more complex computations than the usual multiplication/summation of traditional MLP neurons. Moreover, some of this performance difference might come from how optimized the implementation is. It would be an interesting next step to study the relation of size to performance for KANs and MLPs, since, intuitively, the expressivity of splines should allow for smaller networks.
Table 4: Training time per epoch on NCI1 dataset (i. e., only the forward and backward pass, looped over all batches, averaged over $50$ epochs). All compared models consist of $5$ message passing layers, and of $2$ hidden layers in their respective MLP/KAN. Batch size is set to $128$.
Architecture
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 16 NA NA 3,522 0.368
GIN 64 NA NA 44,802 0.372
GIN 256 NA NA 670,722 0.433
GIN 512 NA NA 2,652,162 0.742
KAGIN 16 1 1 6,752 0.474
KAGIN 32 1 1 21,696 0.475
KAGIN 64 1 1 76,160 0.475
KAGIN 256 1 1 1,091,072 0.732
KAGIN 16 1 3 10,048 0.658
KAGIN 32 1 3 32,384 0.659
KAGIN 64 1 3 113,920 0.792
KAGIN 256 1 3 1,635,328 2.159
KAGIN 16 4 1 11,696 0.476
KAGIN 32 4 1 37,728 0.478
KAGIN 64 4 1 132,800 0.485
KAGIN 256 4 1 1,907,456 1.074
KAGIN 16 4 3 14,992 0.662
KAGIN 32 4 3 48,416 0.726
KAGIN 64 4 3 170,560 1.005
KAGIN 256 4 3 2,451,712 3.092
4.3 Graph Regression
Datasets.
We experiment with two molecular datasets: (1) ZINC-12K [18], and (2) QM9 [32]. ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important chemical property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen (H), Carbon (C), Oxygen (O), Nitrogen (N), and Flourine (F) atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation and a test set according to a $80\%/10\%/10\%$ split.
Experimental setup.
We perform grid search to select values for the different hyperparameters. For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [19]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
Table 5: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on the ZINC-12K and QM9 datasets.
ZINC-12K QM9
GIN 0.4131 $\pm$ 0.0215 0.0969 $\pm$ 0.0017
KAGIN
0.3000
$\pm$ 0.0332
0.0618
$\pm$ 0.0007
Results.
The results are shown in Table 5. We observe that on both considered datasets, KAGIN significantly outperforms the GIN model. Note that these datasets are significantly larger (in terms of number of samples) compared to the graph classification datasets of Table 3. More specifically, KAGIN offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN. Those improvements suggest that KANs might be more effective than MLPs in regression tasks.
5 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. In the classification tasks, there does not appear to be a clear winner, with both architectures outperforming each other on some datasets. In the graph regression task, however, preliminary results seem to indicate that KAN has an advantage over MLP. This paper shows, through its preliminary results, that such KAN-based GNNs are valid alternatives to the traditional MLP-based models. We thus believe that these models deserve the attention of the graph machine-learning community.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove highly relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Finally, a thorough study of the effect of the different hyperparameters could be leveraged, allowing to fully exploit the richness of splines while retaining small networks.
6 Acknowledgements
This work was partially supported by the The Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry.